To: J. Albertine, D. Helms, W. Skinner
cc: Division (M7), FEL Coordination Group
From: F. Dylla
Subject: IRFEL Weekly Report, June 30-July 3, 1997
Date: July 3, 1997
The Martitime Technical Advisory Committee report was received
Feed back from the recent Laser Processing Consortium meeting
has been very positive.
Highlights for the week include the move of the injector drive
laser to the FEL building. The final items for the FEL building
are underway for completion.
Injector drive laser
The drive laser was moved to the FEL building and mounted on the
optical table. We will complete moving all the optics this week.
We have established power to the laser, and today (7/3) are completing
power connections to the chiller system. If the LSOP is signed
today, and certified, we can start the laser to confirm it's still
Regarding the beam-transport system:
The result from the first magnetic measurement of the first optical
chicane dipole from Everson Electric indicates that most of the
manufacturing techniques are right. The transverse scans of
the field at several sections of the gap indicated flat field
within 2 parts per 10,000 at all levels of excitation over a 9
cm. wide path. The specified width for this degree of flatness
throughout the magnet is 7.2 cm.
Machining of the cores and parts for the remainder of the bending
magnets continues. Coils for injection extraction dipoles are
now complete, as are several sets of the coils for reverse bends.
Winding of the 180 degree coils is about to start.
The modifications to the magnet measurement stand for measuring
uniformity within a family are in fabrication with delivery expected
in two weeks and effort is being made to staff up the test stand
for two shift operation and the use of a spare CEBAF power supply
to speed the measurement effort.
The request for quotation for the Panofsky trim quad was sent
out. Bids are due on July 14. Sextupole cores and coils remain
on track for a delivery on July 8
In the corrector area, preliminary prints of the core and coil
for the phasing dipoles and conceptual layouts of the two mu metal
correctors were completed. Steve Lassiter used TOSCA to home
in on a constant perimeter design for the air core corrector horizontal
coil that exhibits less than one percent variation in the X plane
and less than 1.4 % in the Y Plane.
Welding of the bodies for optical chicane chambers started, with
a projected completion date of July 16. The welders are using
the bore scope to inspect for lack of penetration and are rewelding
the non penetrated areas - a good feedback loop that should lead
to more consistent welds the first time. At the same time, a
pump down of the 60 inch test piece chamber, that didn't seem
to be retarded, indicated that contamination from the machining
oil that the X Chamber experienced may not be significant. Our
plan is that further pumping will establish a base pressure by
late next week. At that time we will decide whether rework of
the X and Y chamber bodies is necessary. The back checking of
the drawings for the parts for the reverse bend chambers was completed.
Sign-off is expected early next week. Materials are in at Master
Machine for starting fabrication of the two chambers for the 180
The girders for the injections/extraction dipoles were received.
Drawings for the arc dipole stands and the stands and girders
for the trim quads and sextupoles are in checking. Design of
the beam pipes and stands for the dump lines and shielding and
the short quad triplet after the second arc is now nearing completion.
In the power supply area, the vendor supplying the power supplies
for injection and extraction dipoles, has completion scheduled
for early September. The supplies will have to be sent air freight.
Shunts are complete. The temperature switch chassis and shunt
resistors need assembly.
Regarding cryomodule fabrication:
The cryounit assembly continues to progress with an increase in
production rate. This is relieving some cryomodule completion
schedule stress. The final cryounit is scheduled for completion
8 July 1997.
Cryomodule assembly continues with the first half of the cryomodule
passing leak check of the primary and secondary cryogenic circuits.
During preparation for the installation of the third cryomodule
beam tube a cavity pair isolation valve was found to leak. This
resulting in a partial venting of the cavity pair to air. The
event occurred while the area was under the cover of filtered
laminar flow air. The plan is to continue with a modified assembly
which will minimize contamination potential for the cavity pair.
Based on cavity pair assembly experience with venting cavities
it is expected that performance of the cavity pair will continue
to be acceptable. The performance will not be measured until the
completed cryomodule is installed in the FEL building. Schedule
is tight but still achievable.
Regarding electron-beam instrumentation and controls (I&C):
The network will be operational; Thursday, July 3! The Injector
Test Stand (ITS) server was brought to the FEL building for operating
the drive laser, track the quarter cryomodule cooldown, and commission
the injector RF. The FEL servers will be moved in the FEL building
next week, and they will be brought online for new equipment.
The ITS driven equipment will be converted to the new IOC names
and server in two to four weeks. At that point, the ITS server
will go back to the ITS. Countless cables for instrumentation
and networks have been installed and terminated. Numerous updates
have been made to the web pages: http://www.jlab.org/accel/fel/documentation/evans/felbair1.html.
This site is updated daily reflecting state of installation and
nailing down of details.
Software requirements continue to be fed to the software group.
The personnel safety system continues to make progress toward
the mid-August operational date. The Machine Protection design
review material was handed out to the review committee, the review
will be held following the July 4th accelerator shutdown.
The vacuum control crates are being wired by the accelerator support
group. The ion pump power supplies are complete and waiting control
cards to begin final testing.
WBS 9.4 IR Demo wiggler
Three long flattened vacuum tubes with the required straightness
and dimensions were fabricated this week. We are waiting for
the welding techniques to be proven before making a full chamber.
The girders for the quadupoles on the wiggler bench were successfully
modified to provide for better relative alignement between the
two quadrupoles on the girder. The wiggler bench is now ready
for delivery to the FEL facility. Alignment targets were designed
for boresighting the quadrupoles and wiggler to verify that their
mechanical centers are aligned along a straight line.
Discussions with the vendor (Spawr) for the optical transport
mirrors confirms that they were polished incorrectly, and must
be returned. We plan to do so today (7/3). The balance of the
order, correctly polished, will be shipped next Monday. The collimator
mirror will also ship at that time. We anticipate receipt of
the collimator today. The optical transport line mounting through
the User Labs is mostly complete. We determined that neither
the LCW nor the HVAC were responsible for the ~ 28Hz vibration
the Mach-Zehnder interferometer detects. The nearest source of
vibration are the motors on the cooling water towers for the CHL,
one seemed to be out of balance. This has been remedied, so we
will check again at the next opportunity. This will be later
next week, as we are resetting the position of the optical tables.
Additional PARMELA simulations of the beam dynamics from the photocathode
to the wiggler with the present "long" gun configuration
(cathode-anode spacer in place) were done, with the result that
an injector setup was found that gave the short bunch length at
the injection point that was desired for using the interferometric
bunch-length monitors being procured from the University of Georgia.
The setup is, however, very sensitive to the magnetic field in
the solenoid attached to the gun. As mentioned in recent weekly
reports, we have been assessing the expected performance of the
interferometric bunch-length monitors for given bunch structure.
At "long" bunch lengths, the power from the device
would be "low" and rapidly changing as a function of
the wavelength. Nevertheless, for conceivable bunch lengths,
there is adequate power to perform relative measurements, i.e.,
to monitor bunch length. Cross calibration to an absolute measurement
made with a back-phasing technique is planned. In short, we presently
are not envisioning any serious difficulty using the interferometric
We developed a detailed strawman of the IRFEL commissioning schedule
up to Christmas. The basic strategy is to finish the machine
prior to beam operations, and also to check out subsystems without
beam as early as possible. A probable scenario is to conduct
beam operations during swing and owl shifts in October, with last-minute
cleanup going on during the day shift. Before Christmas we should
have fully tested the injector and cryomodule operation, and we
could conceivably have a start on the coherent-synchrotron-radiation
experiments, these experiments more likely will start in January.
In conjunction with this planning, we are cross-correlating the
activities with the schedule for developing the supporting software.
We decided on an approach to the problem of aligning the four
quadrupoles on the wiggler girder to tight (100-micron-level)
tolerance. We were able easily to find four quadrupoles out of
the recent shipment of QBs (which are thick-lens focusing magnets)
that have twists less than 1 mrad, permitting use of our CEBAF-standard
alignment procedure to get close to the required tolerance. Specifically,
we can use a borescope to align the mechanical centers to 100-micron
uncertainty. However, an additional uncertainty arises because
we can only measure the magnetic center of each magnet to within
about 100 microns relative to its mechanical center. These uncertainties
add in quadrature, thereby exceeding the budget. Thus, to resolve
the discomfort, our plan is to put all four quads and the wiggler
in place on the wiggler girder, aligning these components as closely
as possible, and then checking the alignment of the magnetic centers
with the beam's reference orbit using a pulsed-wire technique.
Beforehand, we will need to establish the precision of the pulsed-wire
measurement. We also, incidentally, are going to measure the
twists of all of the QB quadrupoles in the shipment.
This was a week of details on the facility. Lots of painting
touch up. Sidewalks went in out front on one side. Parking areas
were graded and prepared for asphalt. The entrance area brickwork
was installed. The HVAC was balanced. Punch lists were prepared
for close out. Jefferson Lab locks went on the doors. Columns
covers were added (fire engine red!). Partitions went up in the
womens bath. Ceiling tiles went up everywhere. The vanity was
finished in the men's bathroom. Shower curtains were hung. The
elevator was inspected but failed on fan, sticky door, and smoke
sensor. Downstairs more stands were grouted. Remaining injector
stands were placed and are ready for survey. Lots of RF racks
powered up and much signal hookup and checkout went on. One of
the 50 kW klystrons is almost completely placed, hooked up, and
connected to water, etc. The photoinjector laser went into the
clean room. It is being hooked up to the chiller. Several problems
with the HVAC in the clean room were dealt with but we suspect
the cooling system is still not up to snuff. The manufacturer
has been called. Networks are being hooked up throughout the
building. Network servers are going in. Phones are going in everywhere.
Air and LCW piping is being welded in upstairs and down. Helium
gas has been circulating in the cryogenic pipes. We are ready
to turn on LHe next week.